# Physics-informed deep learning for infectious disease forecasting

### v1

@@ Abstract
Accurate forecasting of contagious illnesses has become increasingly important to public health policymaking, and better prediction could prevent the loss of millions of lives. To better prepare for future pandemics, it is essential to improve forecasting methods and capabilities. In this work, we propose a new infectious disease forecasting model based on physics-informed neural networks (PINNs), an emerging area of scientific machine learning. The proposed PINN model incorporates dynamical systems representations of disease transmission into the loss function, thereby assimilating epidemiological theory and data using neural networks (NNs). Our approach is designed to prevent model overfitting, which often occurs when training deep learning models with observation data alone. In addition, we employ an additional sub-network to account for mobility, vaccination, and other covariates that influence the transmission rate, a key parameter in the compartment model. To demonstrate the capability of the proposed model, we examine the performance of the model using state-level COVID-19 data in California. Our simulation results show that predictions of PINN model on the number of cases, deaths, and hospitalizations are consistent with existing benchmarks. In particular, the PINN model outperforms the basic NN model and naive baseline forecast. We also show that the performance of the PINN model is comparable to a sophisticated Gaussian infection state space with time dependence (GISST) forecasting model that integrates the compartment model with a data observation model and a regression model for inferring parameters in the compartment model. Nonetheless, the PINN model offers a simpler structure and is easier to implement. In summary, our results show that the proposed forecaster could potentially serve as a new computational tool to enhance the current capacity of infectious disease forecasting.
@@ 1 Introduction
COVID-19 was the third largest cause of mortality in the United States in 2022 [1] and, in addition to Ebola, Chikungunya, Zika, and mPox, was one of five global outbreaks in the past decade caused by emerging pathogens. Climate change is expected to exacerbate disease transmission risks, which before COVID-19 already accounted for roughly 25% of global mortality [2, 3]. In the midst of an exponentially growing outbreak, public health officials must respond rapidly in the face of intense uncertainty [4]. Recent developments in infectious disease forecasting have sought to improve our ability to anticipate future epidemiological trends, such as the number of reported cases, hospitalizations, or deaths from a disease, and are increasingly used during outbreaks. Forecasting provides a concrete way to integrate data with epidemiological knowledge and can be used to inform the allocation of critical resources such as antiviral drugs or ventilators, the implementation of non-pharmaceutical interventions, and the design of vaccine trials [4, 5].
@@
Over the past decade, alternative approaches to tackling complex social problems have emerged, including artificial intelligence (AI) and machine learning (ML). These approaches have the advantage of efficiently identifying patterns in data and can, in principle, accommodate multiple disparate data streams. However, conventional AI/ML methods require a large amount of training and labeled data to ensure good performance [24, 25, 26, 27]. Furthermore, AI/ML models may provide unrealistic predictions, especially for non-stationary systems, as they are not constrained by the mechanisms driving transmission. Furthermore, data-driven models have historically struggled with predicting epidemics because they often fail to differentiate real trends from noise in the data collecting process [28]. Recently, AI-based solvers for partial differential equations and ordinary differential equations (PDEs and ODEs) have attracted growing attention due to their ability to assimilate the underlying scientific laws represented by ODEs or PDEs with the learning properties of neural networks (NNs). Combining AI with dynamical systems models can help to infer unknown model parameters as constants or functions using limited data. This enables learning from “small data” as we explicitly utilize the constraints from the physical or biological laws [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Physics-informed neural networks (PINNs) [42] is such a framework fusing data and compartment models to make predictions through the loss function. As illustrated in Fig. 1, the left side of the network represents the fully connected neural network (FNN) (physics-uninformed part), while the right side describes the physics-informed part. The loss function considers the contributions from both the data and compartment models. PINNs have been used to solve scientific and engineering problems, including both forward problems and inverse problems [43, 44, 45, 46, 47, 48, 49].
Figure 1: Structure of the proposed PINNs model for infectious disease forecasting. The PINNs model includes two sub-networks: one is designed for predicting the state variables in the compartmental model, while the other provides predictions for the time-dependent model parameters. The predictions of the two sub-networks u1 and u2 are directly used to compute the data loss, $L_{data}$, which is a weighted sum of contribution from different state variables. These predictions of neural networks are also plugged into the ODE model for computing the ODE loss $L_{r}$.
Here, we introduce a new disease forecasting model based on PINNs. The compartmental model we used contains nine state variables, as listed in Table 1. In addition to the FNN used in the traditional PINNs model, we employ an additional sub-network to account for mobility, vaccination, and other potential covariates that might affect the transmission rate, a key parameter in the compartment model. To demonstrate the capability of this approach, we evaluate the performance of the model by comparing 1-, 2-, 3- and 4-weeks-ahead forecasts of COVID-19 cases, hospital admissions, and deaths in the state of California using reports of COVID-19, Google’s mobility reports, and vaccination data available each week. We assessed model forecasts using weighted interval score [50], which is a known “proper” score for forecasting. We compared the score of our model with the original compartmental model, a naive baseline forecast, and FNNs trained with data only.
@@ 2.1.1 Input Data
To facilitate a comparison of our model’s forecasting performance with that of various other models, we used the forecasting targets defined by the COVID-19 Forecast Hub [51, 21], including COVID-19 case and death data sourced from the COVID-19 Data Repository available at https://github.com/CSSEGISandData/COVID-19, which is maintained by the Center for Systems Science and Engineering (CSSE) at Johns Hopkins University, Baltimore, MD, USA (JHU)[52]. Furthermore, we used information on hospital admissions from the COVID-19 Reported Patient Impact and Hospital Capacity by State Timeseries and COVID-19 Reported Patient Impact and Hospital Capacity by State datasets compiled by the US Department of Health & Human Services (HHS) available on healthdata.gov. Access to these datasets was facilitated through the Delphi Epidata API, as detailed by Reinhart et al. [53]. For our analysis, the time series of hospitalizations is the sum of the fields labeled previous_day_admission_adult_covid_confirmed and previous_day_admission_pediatric_covid_confirmed in the tables provided by HHS.
We used data on mobility and susceptibility as they relate to transmission rate, a significant time-dependent factor for epidemiology forecasting. In particular, mobility can be measured by the amount of time individuals spend in residential areas, as quantified in Google’s community mobility reports [54]. The details of the calculation of our mobility covariate are as follows. For a given forecast date, we obtained the latest snapshot of Google’s Global_Mobility_Report.csv file on http://web.archive.org that was made before the beginning forecast date in the UTC time zone.
Susceptibility is related to the number of vaccine doses administered in a state. The modeling assumption is that the average susceptibility of previously uninfected persons declines with the number of doses administered. We model an effect on average susceptibility rather than the number of susceptibles because a single dose of a COVID-19 vaccine generally does not fully immunize an individual. We sourced our time series of the number of doses administered from the Git repository at www.github.com/govex/COVID-19, which was created by Johns Hopkins University. These data came from either the states’ public dashboards or the Centers for Disease Control and Prevention Vaccine Tracker. Since the reported numbers are cumulative, the creators of this dataset used the larger values from those two data sources each day as the most up-to-date value. To obtain a covariate free of missing values, we assumed that all values before the first reported value were zero. Missing values within the time series were imputed with linear interpolation.
@@ 2.1.2 Data preprocessing
The time series of cases, deaths, hospitalizations, mobility, and cumulative vaccine doses were provided at a daily resolution, but our prediction targets are weekly totals for cases, deaths, and hospitalizations. To be consistent with the targets, we preprocessed these data, including cases, deaths, hospitalizations, mobility, and cumulative vaccine doses, by calculating a 7-day average and transforming the daily dataset to a weekly time series. This smoothing strategy also reduces fluctuations in the original dataset. The hospital admissions time series is different from other datasets as it is recorded later than the cases and deaths time series, and it did not become available until July 2020. At the beginning of the hospital admissions time series, a misleading trend also arose because the number of hospitals providing data was increasing. To align the targets, we discarded data from the first 20 weeks. Additionally, near the end of the time series, there is a small and incomplete peak after two major peaks. We choose to model only the two major peaks and discard the small incomplete peak after week 110. The two major peaks are driven by different virus variants (Beta and Delta) with changing transmission rates, which is sufficient to evaluate the performance of our model.
Fig. 2 shows the original (upper panel) and preprocessed (lower panel) data. In addition, since the time series differ greatly in scale, we perform linear scaling to transform them to a similar scale around 0 to 1, as shown in the lower panel of Fig. 2.
@@ 2.2 Physics-informed neural networks
Table 1: State variables in the employed compartment model by O’Dea and Drake [21].
Physics-informed neural networks (PINNs), which integrate traditional compartment models and data using NNs, have emerged as one of the most influential models in scientific machine learning. We integrate the epidemiology model proposed in [21], which is a compartmental model for COVID-19 with time-dependent parameters, into the proposed PINNs model. The ODE system contains nine equations with nine state variables as well as 11 parameters. The definitions for each state variable are shown in Table 1. Three of the nine state variables ($Z_{r}$, $D_{r}$ and $A$) are observable and were used as targets for training and evaluation. The other six state variables are unobservable and implicit.
@@
Table 2: Model parameters used in the compartment model by O’Dea and Drake [21].
The ODEs are summarized in the physics-informed part in Fig. 1. The definitions and values for each parameter are shown in Table 2. For comparison, we generally adopt the model parameters following the work of [21]: 7 are considered to be known and constant, while three parameters were estimated. Transmission rate $\beta_{t}$ is considered a critical time-dependent parameter and was predicted by a sub-network as shown in Fig. 1. The parameters $p_{h}$ and $p_{d}$ are assumed to be constant and were estimated during model training.
As depicted in Fig. 1, the NNs in PINNs consist of two sub-networks. Both of the sub-networks take time as inputs. Each sub-network consists of three hidden layers with 50 neurons in each layer. One sub-network outputs nine state variables corresponding to the nine variables in the ODE system. This sub-network aims to fit three targets, namely cases, deaths, and hospitalizations, and all nine outputs were used to calculate the loss on the ODEs. The second sub-network outputs three variables, namely mobility, cumulative vaccine doses, and transmission rate. This sub-network aims to fit the mobility data and cumulative vaccine doses time series and outputs transmission rate, which was also used to calculate the loss on the ODEs.
The constraint of the compartmental theory was incorporated into the PINNs training process by adding the loss on the ODEs to the loss on the data, i.e. the total loss was given by the weighted sum of data loss and ODE loss:
$\mathcal{L}(\theta)=w_{data}\mathcal{L}_{data}(\theta)+w_{ODE}\mathcal{L}_{ODE}(\theta)$ (1)
where $\theta$ denotes the PINNs parameters, $\mathcal{L}_{data}$ denotes data loss, $\mathcal{L}_{ODE}$ denotes ODE loss, $w_{data}$ denotes the weight for data loss, and $w_{ODE}$ denotes the weight for ODE loss.
We denote the PINNs as $\bm{y}_{\theta}(\bm{t})$ where $\bm{t}$ denotes input time and $\bm{y}=[y_{1},y_{2},y_{3},y_{4},y_{5}]^{T}$ denotes output with $y_{1}$ to $y_{5}$ representing cases, deaths, hospitalizations, mobility and cumulative vaccine doses respectively. We used the mean squared error (MSE) as the metric for training with the following loss function for data:
@@
where $N_{data}$ is the number of training data and $\{\bm{t}_{i},\bm{y}_{i}\}$ represents the $i$th data point. Then, we trained the PINNs $\bm{y}_{\theta}$ by minimizing the loss function $\mathcal{L}(\theta)$. The Adam optimizer [55] was used with $1\times 10^{-3}$ learning rate for $100,000$ epochs. The hyperparameters, including the learning rate and the weight coefficients, are manually tuned to optimize the prediction accuracy. We trained and tested on a GPU named NVIDIA GeForce RTX 3090.
ODE loss was defined as the sum of the loss on the initial conditions and the loss on residuals.
$\mathcal{L}_{ODE}(\theta)=w_{0}\mathcal{L}_{0}(\theta)+w_{r}\mathcal{L}_{r}(\theta)$ (3)
where $\mathcal{L}_{0}$ denotes initial condition loss, $\mathcal{L}_{r}$ denotes residual loss, $w_{0}$ denotes the weight for initial condition loss, and $w_{r}$ denotes the weight for residual loss.
@@
We trained PINNs using historical data from July 2020 to April 2022. The model was developed with a rolling window approach, allowing continuous predictions for the following 1 to 4 weeks. Initially, we trained PINNs on preprocessed data spanning 1 to 17 weeks and made predictions for weeks 18 to 21. We then shifted the training window by one week, training new PINNs with data from weeks 1 to 18 and predicting for weeks 19 to 22. This process was repeated until we trained PINNs using data from weeks 1 to 89 and predicted for weeks 90 to 93. Consequently, each week from 21 to 90 had been treated as the subsequent 1, 2, 3, and 4 weeks during the rolling window approach.

### v2

@@ Abstract
Accurate forecasting of contagious illnesses has become increasingly important to public health policymaking, and better prediction could prevent the loss of millions of lives. To better prepare for future pandemics, it is essential to improve forecasting methods and capabilities. In this work, we propose a new infectious disease forecasting model based on physics-informed neural networks (PINNs), an emerging area of scientific machine learning. The proposed PINN model incorporates dynamical systems representations of disease transmission into the loss function, thereby assimilating epidemiological theory and data using neural networks (NNs). Our approach is designed to prevent model overfitting, which often occurs when training deep learning models with observation data alone. In addition, we employ an additional sub-network to account for mobility, cumulative vaccine doses, and other covariates that influence the transmission rate, a key parameter in the compartmental model. To demonstrate the capability of the proposed model, we examine the performance of the model using state-level COVID-19 data in California. Our simulation results show that predictions of PINN model on the number of cases, deaths, and hospitalizations are consistent with existing benchmarks. In particular, the PINN model outperforms naive baseline forecasts and various sequence deep learning models, such as Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, Gated Recurrent Units (GRUs), and Transformer models. We also show that the performance of the PINN model is comparable to a sophisticated Gaussian infection state forecasting model that combines the compartmental model, a data observation model and a regression model for inferring parameters in the compartmental model. Nonetheless, the PINN model offers a simpler structure and is easier to implement. In summary, we perform a systematic study of the predictive capability of the PINN model in forecasting the dynamics of infectious diseases and our results showcase the potential of the proposed model as an efficient computational tool to enhance the current capacity of infectious disease forecasting.
@@ 1 Introduction
COVID-19 was the third largest cause of mortality in the United States in 2022 [1] and, in addition to Ebola, Chikungunya, Zika, and mPox, was one of five global outbreaks in the past decade caused by emerging pathogens. Climate change is expected to exacerbate disease transmission risks, which before COVID-19 already accounted for roughly 25% of global mortality [2, 3]. In the midst of an exponentially growing outbreak, public health officials must respond rapidly in the face of intense uncertainty [4]. Recent developments in infectious disease forecasting have sought to improve our ability to anticipate future epidemiological trends, such as the number of reported cases, deaths, or hospitalizations from a disease, and are increasingly used during outbreaks. Forecasting provides a concrete way to integrate data with epidemiological knowledge and can be used to inform the allocation of critical resources such as antiviral drugs or ventilators, the implementation of non-pharmaceutical interventions, and the design of vaccine trials [4, 5, 6].
@@
Over the past decade, alternative approaches to tackling complex social problems have emerged, including artificial intelligence (AI) and machine learning (ML). These approaches have the advantage of efficiently identifying patterns in data and can, in principle, accommodate multiple disparate data streams. However, conventional AI/ML methods require a large amount of training and labeled data to ensure good performance [25, 26, 27, 28]. Furthermore, AI/ML models may provide unrealistic predictions, especially for non-stationary systems, as they are not constrained by the mechanisms driving transmission. Furthermore, data-driven models have historically struggled with predicting epidemics because they often fail to differentiate real trends from noise in the data collecting process [29]. Recently, AI-based solvers for partial differential equations (PDEs) and ordinary differential equations (ODEs) have attracted growing attention due to their ability to assimilate the underlying scientific laws represented by ODEs or PDEs with the learning properties of neural networks (NNs). Combining AI with dynamical systems models can help to infer unknown model parameters as constants or functions using limited data. This enables learning from “small data” as we explicitly utilize the constraints from the physical or biological laws [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Physics-informed neural networks (PINNs) [44] provide a powerful framework that integrates data with compartmental models by incorporating both into the loss function. As illustrated in Fig.1A, the left side of the architecture represents a fully connected neural network (FCNN), corresponding to the physics-uninformed component, while the right side captures the physics-informed component. The loss function is designed to account for contributions from both observed data and the underlying compartmental models. PINNs have been successfully applied to a range of scientific and engineering challenges, encompassing both forward and inverse problems [45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
Figure 1: A. Schematic of the proposed PINNs model for infectious disease forecasting. The model comprises two sub-networks: the upper sub-network predicts the state variables in the compartmental model, while the lower sub-network estimates the time-dependent model parameters. The output $u_{1}$ represents the nine compartmental state variables, including $X$, $L$, $Z$, $Z_{r}$, $H$, $A$, $D$, $D_{r}$. The output $u_{2}$ represents factors including mobility, cumulative vaccine doses, and transmission rate. The data loss $L_{data}$ is a weighted sum of observable state variables and factors. B. Schematic of the rolling window approach. As training data accumulates over time, the PINNs model continuously updates and generates forecasts for the subsequent 1–4 weeks.
Here, we introduce a novel disease forecasting model based on PINNs. The compartmental model includes nine state variables (Tab. 1), with the PINNs architecture comprising two sub-networks that take time $t$ as input. The upper sub-network predicts all nine state variables, contributing to the ODE loss $\mathcal{L}_{\text{ODE}}$, while three observable variables – reported cases ($Z_{r}$), hospitalizations ($A$), and reported deaths ($D_{r}$) – also contribute to the data loss $\mathcal{L}_{\text{data}}$. The lower sub-network captures the relationship between transmission rate and two associated factors: mobility and cumulative vaccine doses. It outputs predictions for all three, with transmission rate feeding into $\mathcal{L}_{\text{ODE}}$, and the other two informing $\mathcal{L}_{\text{data}}$. This design allows for seamless integration of mechanistic modeling with observed data, enhancing forecasting performance.
To demonstrate the capability of this approach, we evaluated the performance of the model by comparing 1-, 2-, 3- and 4-weeks-ahead forecasts of COVID-19 cases, deaths, and hospitalizations in the state of California using reports of COVID-19, Google’s mobility reports, and vaccination data available each week. We assessed model accuracy on both point forecasting and quantile forecasting [55]. We compared the score of our model with a naive baseline forecast, various sequence deep learning models, such as recurrent neural networks (RNNs), long short-term memory (LSTM) networks, gated recurrent units (GRUs), and Transformer models as well as a traditional mathematical model.
@@ 2.1.1 Input Data
To facilitate a comparison of our model’s forecasting performance with that of various other models, we used the forecasting targets defined by the COVID-19 Forecast Hub [56, 22], including COVID-19 case and death data sourced from the COVID-19 Data Repository available at https://github.com/CSSEGISandData/COVID-19, which is maintained by the Center for Systems Science and Engineering (CSSE) at Johns Hopkins University, Baltimore, MD, USA (JHU) [57]. Furthermore, we used information on hospitalizations from the COVID-19 Reported Patient Impact and Hospital Capacity by State Timeseries and COVID-19 Reported Patient Impact and Hospital Capacity by State datasets compiled by the US Department of Health & Human Services (HHS) available on healthdata.gov. Access to these datasets was facilitated through the Delphi Epidata API, as detailed by Reinhart et al. [58]. For our analysis, the time series of hospitalizations is the sum of the fields labeled previous_day_admission_adult_covid_confirmed and previous_day_admission_pediatric_covid_confirmed in the tables provided by HHS.
We used data on mobility and cumulative vaccine doses due to their association with the transmission rate, a critical time-dependent parameter in epidemiological forecasting. In particular, mobility can be measured by the amount of time individuals spend in residential areas, as quantified in Google’s community mobility reports [59]. For a given forecast date, we obtained the latest snapshot of Google’s Global_Mobility_Report.csv file on http://web.archive.org that was made before the beginning forecast date in the UTC time zone.
Cumulative vaccine doses represent the total number of vaccine doses administered within a given state. Our modeling assumes that the average susceptibility of individuals who have not been previously infected decreases as the number of administered doses increases. Rather than modeling the number of susceptible individuals directly, we model the effect of vaccination on average susceptibility, acknowledging that a single dose of a COVID-19 vaccine typically does not confer full immunity. We obtained the time series data on administered vaccine doses from the GitHub repository maintained by Johns Hopkins University (www.github.com/govex/COVID-19). This dataset compiles information from state public dashboards and the Centers for Disease Control and Prevention (CDC) Vaccine Tracker. For each day, the dataset reports the higher value between the two sources to provide the most current estimate. To generate a covariate without missing values, we assumed all values prior to the first reported entry were zero, and any missing values within the series were imputed using linear interpolation.
@@ 2.1.2 Data preprocessing
The time series data for cases, deaths, hospitalizations, mobility, and cumulative vaccine doses were originally provided at a daily resolution, whereas our prediction targets are weekly totals for cases, deaths, and hospitalizations. To ensure consistency with the prediction targets, we preprocessed these datasets by computing a 7-day moving average and aggregating the daily values into a weekly time series. This smoothing strategy not only aligns the temporal resolution with the forecasting targets but also reduces short-term fluctuations in the original data.
The hospitalization data differ from the other time series in several key aspects. First, the hospitalization records were made available later than the case and death data, with reliable reporting beginning in July 2020. Second, the initial portion of the hospitalization time series is affected by an artificial upward trend caused by the gradual increase in the number of hospitals reporting data. To mitigate this bias, we excluded the first 20 weeks of hospitalization data from our analysis.
Toward the end of the time series, a small, incomplete peak appears following two major peaks. To maintain focus on well-characterized dynamics, we chose to model only the two major peaks and discarded data after week 110. These two peaks correspond to different viral variants (Beta and Delta), each associated with distinct transmission dynamics, and are sufficient for evaluating the performance of our forecasting model.
Fig. 2 illustrates both the original (upper panel) and preprocessed (lower panel) time series data. To address the large differences in scale across variables, we applied linear scaling to normalize each time series to a similar range, approximately between 0 and 1. This normalization is also reflected in the lower panel of Fig. 2, facilitating better visualization and more stable model training.
@@ 2.2 Physics-informed neural networks
Table 1: State variables in the employed compartmental model by O’Dea and Drake [22].
Physics-informed neural networks (PINNs), which combine traditional compartmental models with observational data through neural networks, have emerged as one of the most influential approaches in scientific machine learning. In this work, we incorporate the epidemiological model proposed in [22] – a compartmental model for COVID-19 with time-dependent parameters – into our PINNs framework. This model is governed by a system of nine ODEs, involving nine state variables and eleven parameters. The definitions of the state variables are provided in Tab. 1. Among these, three state variables – reported cases $Z_{r}$, reported deaths $D_{r}$, and hospitalizations $A$ – are observable and used as targets for model training and evaluation. The remaining six state variables are unobservable and are treated as latent variables within the PINNs framework.
@@
Table 2: Model parameters used in the compartmental model by O’Dea and Drake [22].
The system of ODEs is incorporated into the physics-informed component of the model, as illustrated in Fig. 1A. Definitions and values for each model parameter are provided in Tab. 2. For consistency and comparability, we generally adopt the parameter settings from [22]: seven parameters are treated as known and constant, while three are estimated during training. Among them, the transmission rate $\beta_{t}$ is considered a critical time-dependent parameter and is predicted by the lower sub-network, as depicted in Fig. 1A. The remaining two parameters, $p_{h}$ and $p_{d}$, are assumed to be constant and are integrated into the model parameters to be estimated during the training process.
As depicted in Fig. 1A, the PINNs architecture comprises two sub-networks, both taking time $t$ as input. Each sub-network consists of three hidden layers with 50 neurons in each layer. The upper sub-network predicts the values of the nine state variables, all of which are used to compute the ODE loss, denoted as $\mathcal{L}_{ODE}$. Among these, three variables – reported cases $Z_{r}$, hospitalizations $A$, and reported deaths $D_{r}$ – are observable and contribute to the data loss $\mathcal{L}_{data}$.
The lower sub-network is designed to implicitly capture the relationship between the transmission rate and two closely associated factors: mobility and cumulative vaccine doses. It outputs predictions for mobility, cumulative vaccine doses, and the transmission rate. The transmission rate is a key time-dependent parameter in the compartmental model and is involved in the computation of $\mathcal{L}_{ODE}$, while mobility and cumulative vaccine doses has corresponding reported data and contribute to $\mathcal{L}_{data}$. This PINNs framework enables the integration of mechanistic modeling with real-world data for improved forecasting accuracy.
The constraint imposed by the compartmental model was integrated into the PINNs training process by combining the ODE loss with the data loss. Specifically, the total loss is defined as a weighted sum of the data loss and the ODE loss:
$\mathcal{L}(\theta)=\mathcal{L}_{data}(\theta)+w_{ODE}\mathcal{L}_{ODE}(\theta)$ (1)
where $\theta$ denotes the PINNs parameters, $\mathcal{L}_{data}$ represents data loss, $\mathcal{L}_{ODE}$ represents ODE loss, and $w_{ODE}$ represents the weight for ODE loss.
We denote the PINNs model as $\bm{y}_{\theta}(\bm{t})$, where $\bm{t}$ represents the input time, and $\bm{y}=[y_{1},y_{2},y_{3},y_{4},y_{5}]^{T}$ denotes the output vector, with $y_{1}$ to $y_{5}$ corresponding to cases, deaths, hospitalizations, mobility, and cumulative vaccine doses, respectively. We used the mean squared error (MSE) as the training metric, with the following loss function for the data:
@@
where $N_{\text{data}}$ denotes the number of training data points, and $\{\bm{t}_{i},\bm{y}_{i}\}$ represents the $i$th data pair. The PINNs $\bm{y}_{\theta}$ were trained by minimizing the loss function $\mathcal{L}(\theta)$.
The ODE loss was defined as the sum of the loss on the initial conditions and the loss on the residuals.
$\mathcal{L}_{ODE}(\theta)=\mathcal{L}_{0}(\theta)+\mathcal{L}_{r}(\theta)$ (3)
where $\mathcal{L}_{0}$ represents initial condition loss, $\mathcal{L}_{r}$ represents residual loss.
@@
Overfitting is a common challenge in time series forecasting. To mitigate this, we employ L2 regularization, a widely used technique that adds a penalty term – the sum of the squared network parameters – to the original loss function during training. This regularization discourages overly complex models by penalizing large parameter values. In practice, L2 regularization is well-integrated in PyTorch and can be easily applied by setting the weight_decay parameter in the optimizer.
We used the Adam optimizer [60] with a learning rate of $1\times 10^{-3}$ for 50,000 epochs. L2 regularization with a coefficient of $1\times 10^{-5}$ was applied to reduce overfitting. To balance the scales between $\mathcal{L}_{\text{data}}$ and $\mathcal{L}_{\text{ODE}}$, we set the weight $w_{\text{ODE}}=0.1$. All hyperparameters were manually tuned to optimize predictive accuracy. Training and testing were conducted on an NVIDIA GeForce RTX 3090 GPU.
We trained the PINNs model using historical data from July 2020 to April 2022, employing a rolling window approach, as illustrated in Fig. 1B, to enable continuous predictions for the subsequent 1 to 4 weeks. Initially, PINNs were trained on data from weeks 1 to 17, with predictions made for weeks 18 to 21. The training window was then shifted by one week, retraining the model on data from weeks 1 to 18 and forecasting weeks 19 to 22. This process was repeated until PINNs were trained using data from weeks 1 to 89 and predictions were made for weeks 90 to 93. As a result, each week from 21 to 90 was treated as the subsequent 1-, 2-, 3-, and 4-week forecast in the rolling window procedure.